inverse of y=4(x-2)+9
I can do easy questions but not this please help and quickly ,thanking you in advance please do quickly as in for early tomorrow mourning many thanks

Answer by Edwin McCravy(20054)

inverse of y = 4(x - 2) + 9

I can do easy questions but not
this please help and quickly ,
thanking you in  advance please
do quickly as in for early tomorrow
mourning many thanks 

To find the inverse:

     y = 4(x - 2) + 9

Swap x and y

     x = 4(y - 2) + 9

Solve for y:

     x = 4y - 8 + 9

     x = 4y + 1

 x - 1 = 4y

Divide both sides by 4

 x - 1     4y
------- = ----
   4        4

 x - 1    
------- = y
   4       

     x - 1    
y = ------- 
       4

That's it.

Edwin